The morning was quite short today because most of the crew worked late at night yesterday.
The astronomers went to observe until dawn and the geologist analyzed his samples under UV light.
The astronomy session was quite productive and brought back a handful of pictures. The spectrum of
the Moon could also be recorded and improvements ave been made for the detection of the spectrum of
The UV analysis also wen well and will be concluded by tomorrow for this set of samples.
The astronomy pictures were processed during the day and the geologist finalized the mapping of our
About our other experiments, the balloon project was advanced as well as it is now ready for the
launch scheduled tomorrow.
We experienced an Internet connection problem between 3.00 pm and 7.03pm
Crew Physical Health:
All are well.
There was no EVA today.
No internet connection.
We have access to HALpr and the local network. But the internet
(satellite) connection is
unreachable. Pinging yields 100% lost packages.
Close-range EVA for the launch of the helium balloon.
Long-range EVA: Mineral sampling and measuring streambeds.
If the sky allows it, spectroscopy of sky objects.
(Copy-paste from yesterday, since we didn't get an answer)
1) Could you ack you received our video and the corresponding notification
Also, were you able to read the video file and could you hear the soundtrack?
If you wish so, we can also upload it in full definition (400 Megabytes
Could you tell us what the best time is to do so, to have the best
2a) Would it be possible to add a link to our website on
Its URL is www.mdrs-ucl.be
2b) There is a small error on our biography page:
Thibaud's family name is "Herriau" (and not hArriau).